A school jazz ensemble collected $670 for 300 tickets sold for a concert. If adult tickets were $9 each, and students $4 each, how many of each were sold? Set up equations and solve.
Answer by rfer(12646)

a+s=300 a=300-s
9a+4s=670
9(300-s)+4s=670
2700-9s+4s=670
-5s=-2030
s=406 students
a=670-406
a=264 adults